OpenStudy (anonymous):

Which sequence can be represented by the expression 12n? 144, 132, 120, 108,… 1, 12, 24, 36,… 12, 24, 36, 48,… 6, 18, 30, 42,…

OpenStudy (anonymous):

3 rd one

OpenStudy (anonymous):

Times

OpenStudy (anonymous):

and can ui please explain

OpenStudy (anonymous):

12*1=12 12*2=24 12*3=36 so it's the third one

OpenStudy (anonymous):

@zzr0ck3r makes no difference the answer is basically the table of 12

OpenStudy (anonymous):

thx guys\

OpenStudy (zzr0ck3r):

of course it makes a difference if n runs through the integers, then 0 would need to be there and -12...

OpenStudy (zzr0ck3r):

ahh they defined it as a sequence, so its natural

OpenStudy (anonymous):

@tommyblythe56 This is very simple see the expression 12n which means that 12 multiplied with n so n can take any value from 1 to infinity which basically gives you the multiplication table of 12